OpenStudy (anonymous):

Tough Integral Integrate [(c+x)/(c-x)]^.5 from -c to c

OpenStudy (anonymous):

\[\int\limits_{-c}^{c}\sqrt{\frac{ c+x }{c-x }}dx\]

OpenStudy (anonymous):

\[\int\limits_{-c}^{c} \sqrt{(\frac{c+x}{c-x})}dx\] try \[x=c \times \cos(2t)\]\[dx=-2csin(2t)dt\]

OpenStudy (anonymous):

I will Thanks

OpenStudy (anonymous):

So it's been a really long time since I've done substitution. How do I change the constants of integration?

OpenStudy (anonymous):

What do you mean?This a definite integral so you wouldn't have any constant of integration

OpenStudy (anonymous):

Sorry I mean the interval\[\int\limits_{-c}^{c} \rightarrow \int\limits_{?}^{?}\]

OpenStudy (michele_laino):

Please try this substitution: \[\frac{{c + x}}{{c - x}} = {t^2}\] where t is the new variable of integration

OpenStudy (anonymous):

sorry, to change the limits put \[x=c\] and \[x=-c\] find the value in terms of t for the corresponding limit

OpenStudy (anonymous):

what would dt be?

OpenStudy (anonymous):

This is for a fracture mechanics question. My professor said he substituted b=x/c and db=dx/c and it worked out great but I cant seem to get this